# The Mutual-Visibility Problem In Directed Graphs

### v1

@@ Abstract
The mutual-visibility problem, originally defined for undirected graphs, asks for the size of the maximum set of vertices $S$ such that every pair of vertices in $S$ is connected by a shortest path passing only through vertices in $V\setminus S$. In this paper, we extend this concept to directed graphs, establishing fundamental results for several graph classes. We prove that for Directed Acyclic Graphs (DAGs), the mutual-visibility number $\mu(D)$ is always 1, and for directed cycles of length $n\geq 3$, it is strictly 2. In contrast, we demonstrate that tournaments can support arbitrarily large mutual-visibility sets; specifically, using properties of Paley tournaments, we show that $\mu(T)$ grows linearly with the size of the tournament. On the algorithmic side, we show that while verifying a candidate set is polynomial-time solvable ($O(|S|(|V|+|A|))$), the problem of determining $\mu(D)$ is NP-hard for general digraphs. We also analyze the impact of strong bridges and strongly connected components on the upper bounds of $\mu(D)$.
@@ 1 Introduction
The mutual‐visibility problem on graphs asks for a largest set of vertices such that any two in the set are “mutually visible,” meaning there is at least one shortest path between them whose internal vertices lie outside the set. Formally, a set $S\subseteq V(G)$ is a mutual‐visibility set if for every $\in S, xyx\neq y$, there exists a shortest $x$,$y$-path in $G$ with no internal vertex in $S$. The size of a maximum mutual‐visibility set is the mutual‐visibility number $\mu(G)$. It was first defined in DISTEFANO2022126850 by Di Stefano.
Di Stefano’s seminal work DISTEFANO2022126850 explicitly motivates further exploration into the mutual-visibility problem, particularly within the context of directed graphs. This specific area currently exhibits a notable scarcity of dedicated research. Our work aims to address this gap by establishing fundamental results concerning mutual-visibility in directed graphs.
@@
The concept of mutual visibility was introduced by Di Stefano DISTEFANO2022126850 in the context of undirected graphs. This work was motivated by problems in robot navigation and communication networks, where entities must maintain lines of sight or communication channels without interference from other entities in the swarm. Di Stefano established that the mutual-visibility number is related to, but distinct from, other metric invariants such as the general position number and the geodetic number.
Following this seminal work, Cicerone et al. CICERONE2023114096 expanded the field by investigating a variety of mutual-visibility problems, considering different constraints on the paths (e.g., all shortest paths vs. at least one shortest path). They provided bounds for grids, orbital graphs, and other specific topologies.
However, the literature has remained focused almost exclusively on undirected graphs. To the best of our knowledge, the problem of mutual visibility in directed graphs has not been addressed. The transition to directed graphs introduces significant structural challenges: distance is no longer symmetric ($d(u,v)\neq d(v,u)$), and the existence of a path does not guarantee a return path. These asymmetries render many undirected techniques inapplicable. For instance, while removing an edge in a cycle simply increases distance in the undirected case, it destroys strong connectivity entirely in the directed case. Our work aims to bridge this gap by defining the problem for digraphs and characterizing $\mu(D)$ for fundamental directed structures.
@@ 2 Mutual-Visibilities and Properties in Directed Graphs
Following the definitions of CICERONE2023114096 , we also define mutual-visibility sets.
@@ 4 Tournaments
Historically, the properties of paths in tournaments have been studied via probabilistic methods. It is a standard result in the theory of random tournaments, discussed by Moon MOON1968 , that for sufficiently large $n$, every pair of vertices has approximately $n/4$ common outgoing neighbors. To provide a concrete lower bound for $\mu(T)$, we utilize Paley tournaments, a class of deterministic tournaments that exhibit these ”random-like” properties (often called quasirandom or doubly regular).
@@ 5 Algorithmic Complexity
We continue by addressing the computational complexity of the mutual-visibility problem in directed graphs. Following the foundational work of Di Stefano DISTEFANO2022126850 , we analyze both the complexity of verifying a candidate set and the hardness of determining the mutual-visibility number $\mu(D)$.
@@ 5.2 Hardness of Finding $\mu(D)$
While verification is efficient, finding the maximum size of such a set is computationally intractable. Di Stefano proved that the mutual-visibility problem is NP-complete for general undirected graphs DISTEFANO2022126850 . We extend this result to directed graphs via a simple reduction.
@@
We proceed by reduction from the Mutual-Visibility problem on undirected graphs, which is known to be NP-complete DISTEFANO2022126850 .

### v2

@@ Abstract
The study of mutual visibility has traditionally focused on undirected graphs, asking for the maximum number of vertices that can communicate via shortest paths without intermediate interference from other set members. In this paper, we extend this concept to directed graphs, establishing fundamental results for several graph classes. We prove that for Directed Acyclic Graphs (DAGs), the mutual-visibility number $\mu(D)$ is always 1, and for directed cycles of length $n\geq 3$, it is strictly 2. In contrast, we demonstrate that tournaments can support arbitrarily large mutual-visibility sets; specifically, using properties of Paley tournaments, we show that $\mu(T)$ grows linearly with the size of the tournament. On the algorithmic side, we show that while verifying a candidate set is polynomial-time solvable ($O(|S|(|V|+|A|))$), the problem of determining $\mu(D)$ is NP-hard for general digraphs. We also analyze the impact of strong bridges and strongly connected components on the upper bounds of $\mu(D)$.
@@ 1 Introduction
The mutual‐visibility problem on graphs asks for a largest set of vertices such that any two in the set are “mutually visible,” meaning there is at least one shortest path between them whose internal vertices lie outside the set. Formally, a set $S\subseteq V(G)$ is a mutual‐visibility set if for every $\in S, xyx\neq y$, there exists a shortest $x$,$y$-path in $G$ with no internal vertex in $S$. The size of a maximum mutual‐visibility set is the mutual‐visibility number $\mu(G)$. It was first defined in Di Stefano [2022] by Di Stefano.
Di Stefano’s seminal work Di Stefano [2022] explicitly motivates further exploration into the mutual-visibility problem, particularly within the context of directed graphs. This specific area currently exhibits a notable scarcity of dedicated research. Our work aims to address this gap by establishing fundamental results concerning mutual-visibility in directed graphs.
@@
The concept of mutual visibility was introduced by Di Stefano Di Stefano [2022] in the context of undirected graphs. This work was motivated by problems in robot navigation and communication networks, where entities must maintain lines of sight or communication channels without interference from other entities in the swarm. Di Stefano established that the mutual-visibility number is related to, but distinct from, other metric invariants such as the general position number and the geodetic number.
Following this seminal work, Cicerone et al. Cicerone et al. [2023] expanded the field by investigating a variety of mutual-visibility problems, considering different constraints on the paths (e.g., all shortest paths vs. at least one shortest path). They provided bounds for grids, orbital graphs, and other specific topologies.
However, the literature has remained focused almost exclusively on undirected graphs. To the best of our knowledge, the problem of mutual visibility in directed graphs has not been addressed. The transition to directed graphs introduces significant structural challenges: distance is no longer symmetric ($d(u,v)\neq d(v,u)$), and the existence of a path does not guarantee a return path. These asymmetries render many undirected techniques inapplicable. For instance, while removing an edge in a cycle simply increases distance in the undirected case, it destroys strong connectivity entirely in the directed case. Our work aims to bridge this gap by defining the problem for digraphs and characterizing $\mu(D)$ for fundamental directed structures. Consider the practical implications of this asymmetry in communication networks. In an undirected model, a link between two nodes implies a bidirectional channel. However, in real-world systems like optical networks with isolators or traffic grids with one-way streets, visibility is not inherently reciprocal. A node $u$ might ”see” $v$, but $v$ may require a completely different, longer route to see $u$.
Finally, we distinguish our contributions from recent parallel developments in the undirected setting. Notably, Roy et al. Roy et al. [2025]recently analyzed the vertex visibility number, linking the visibility of a subset $S$ from a fixed root $x$ to the leaves of a shortest-path tree. While their work provides significant insights into the structure of visibility in undirected graphs, their reliance on shortest-path trees is inherently symmetric. In contrast, our work addresses the strictly directed case, where path asymmetry ($u\to v\nleftrightarrow v\to u$) prevents the formation of standard shortest-path trees, requiring a fundamentally different structural analysis based on strong connectivity and cycle decomposition.
@@ 2 Mutual-Visibilities and Properties in Directed Graphs
Following the definitions of Cicerone et al. [2023], we also define mutual-visibility sets.
@@ 4 Tournaments
Historically, the properties of paths in tournaments have been studied via probabilistic methods. It is a standard result in the theory of random tournaments, discussed by Moon Moon [1968], that for sufficiently large $n$, every pair of vertices has approximately $n/4$ common outgoing neighbors. To provide a concrete lower bound for $\mu(T)$, we utilize Paley tournaments, a class of deterministic tournaments that exhibit these ”random-like” properties (often called quasirandom or doubly regular).
@@
Intuitively, tournaments represent the structural opposite of DAGs regarding visibility. While DAGs enforce a strict ordering that prevents any return paths (yielding $\mu(D)=1$), tournaments are maximally connected in the sense that every pair has a direct relationship. The challenge for mutual visibility here lies in finding a ”triangulated” return path of length 2. Paley tournaments, constructed from finite fields, possess quasirandom properties that ensure a dense distribution of these specific 2-step paths, effectively preventing any small set of vertices from blocking all return routes.
@@ 5 Algorithmic Complexity
We continue by addressing the computational complexity of the mutual-visibility problem in directed graphs. Following the foundational work of Di Stefano Di Stefano [2022], we analyze both the complexity of verifying a candidate set and the hardness of determining the mutual-visibility number $\mu(D)$.
@@ 5.2 Hardness of Finding $\mu(D)$
While verification is efficient, finding the maximum size of such a set is computationally intractable. Di Stefano proved that the mutual-visibility problem is NP-complete for general undirected graphs Di Stefano [2022]. We extend this result to directed graphs via a simple reduction.
@@
To prove hardness, we employ a reduction from the undirected version of the Mutual-Visibility problem, established as NP-complete by Di Stefano Di Stefano [2022].